Bracken Corporation is a small wholesaler of gourmet food products. Data regarding the store's operations follow:

o Sales are budgeted at $462,000 for November, $348,000 for December, and $380,000 for January.
o Collections are expected to be 80% in the month of sale, 17% in the month following the sale, and 3% uncollectible.
o The cost of goods sold is 85% of sales.
o The company would like to maintain ending merchandise inventories equal to 70% of the next month's cost of goods sold. Payment for merchandise is made in the month following the purchase.
o Other monthly expenses to be paid in cash are $26,550.
o Monthly depreciation is $20,400.
o Ignore taxes.

Balance Sheet October 31
Assets
Cash $29,000
Accounts receivable, net of allowance for uncollectible accounts 82,400
Merchandise inventory 217,500
Property, plant and equipment, net of $656,000 accumulated depreciation 1,312,000
Total assets 1,640,900
Liabilities and Stockholders' Equity
Accounts payable $300,000
Common stock 840,000
Retained earnings 500,900
Total liabilities and stockholders' equity $1,640,900


The cost of December merchandise purchases would be:

$314,840

$226,100

$295,800

$392,700

question 2)

Rhett Corporation manufactures and sells dress shirts. Each shirt (unit) requires 5 yards of cloth. Selected data from Rhett's master budget for next quarter are shown below:

April May June
Budgeted sales (in units) 28,100 29,100 41,300
Budgeted production (in units) 35,700 38,400 37,300
Desired ending inventory of cloth (in yards) 3,570 3,500 2,760


How many yards of cloth should Rhett plan on purchasing in May?

191,930 yards

179,930 yards

195,570 yards

195,500 yards
Answer #1

1) December purchase = Cost of goods sold+Ending inventory-Beginning inventory

= (348000*85%)+(380000*85%*70%)-(348000*85%*70%)

December purchase = 314840

So answer is a) $314840

2) Purchase in May = (38400*5)+3500-3570 = 191930

So answer is a) 191930 Yards